A ball is projected vertically upward with a speed of 50 m/s. Find (a) the maximum height, (b) the time to reach the maximum height, (c) the speed at half the maximum height. Take g = 10 m/s2.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. u = 50 m/s, g = –10 m/s2 when moving upward, v = 0 (at highest point). a) S = (V^2-u^2)/2a = (0-〖50〗^2)/(2(-10)) = 125 m maximum height reached = 125 m b) t = (v – u)/a = (0 – 50)/–10 = 5 sec c) s’ = 125/2 = 62.5 m, u = 50 m/s, a = –10 m/s2, v2 – u2 = 2as ⇒ v = √(( u^2+2as)) = √(〖50〗^2+2(-10)(62.5) ) = 35 m/s.